Find the general solution to y'' − y' − 2y = 2e3x

1
Expert's answer
2021-09-22T13:32:23-0400

"k^2-k-2=0"


"k=\frac{1\pm\sqrt{1+8}}{2}"


"k_1=-1,\ k_2=2"


"Y=c_1e^{-x}+c_2e^{2x}"


"\tilde{y}=Ae^{3x}"

"\tilde{y}'=3Ae^{3x}"

"\tilde{y}''=9Ae^{3x}"

"9Ae^{3x}-3Ae^{3x}-2Ae^{3x}=2e^{3x}"

"9A-3A-2A=2"

"A=1\/2"


"\tilde{y}=e^{3x}/2"


"y=Y+\tilde{y}=c_1e^{-x}+c_2e^{2x}+e^{3x}/2"